# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

### v2

@@ 5 Results
| Siemens Myomap | ||
| MBSS-T1 | ||
| Method | ICC3 (on Mean) | ICC3 (on Median) | ||||
| FB | BH | FB+BH | FB | BH | FB+BH | |
| w.o motion correction | ||||||
| Siemens Myomap | ||||||
| MBSS-T1 | ||||||
6 Conclusions
This work presents a self-supervised method for motion-robust cardiac T1 mapping. The method utilizes a self-supervised learning approach combined with a pre-trained segmentation framework, enhanced by incorporating a signal decay model within the network’s architecture and an anatomical constraint. This approach enables the generation of corrected T1 maps across various imaging protocols without requiring protocol-specific data collection or extensive training. The main contribution of this work lies in the novel combination of signal decay and anatomical constraints within the registration process, providing a powerful tool for improving motion correction and expanding the reach of quantitative MRI techniques in clinical practice.
Compared to baseline methods, the method demonstrated quantitative improvements in fit quality, Dice score, and Hausdorff distance. A qualitative evaluation by an expert cardiac radiologist indicated the clinical relevance of the generated T1 maps, showing potential for accurate T1 estimation in patients unable to perform breath-hold during MRI scans. This approach may extend the applicability of quantitative cardiac T1 mapping across a broader patient population, ensuring reliable imaging even under challenging conditions.
In addition to its effectiveness in T1 mapping, the proposed approach holds significant potential for application to other quantitative MRI problems. By integrating signal decay modeling with anatomical constraints, this method offers a robust framework that could be adapted to correct motion and improve accuracy in various quantitative MRI sequences, such as T2 mapping, diffusion MRI, and beyond. Furthermore, the flexible framework could be further enhanced by replacing the backbone with more advanced architectures, such as transformer-based models. This substitution could leverage the global attention mechanisms inherent in transformers, potentially improving the capture of complex anatomical structures and further enhancing the accuracy of the generated maps.

### v3

@@ 5 Results
| Voxelmorph | ||
| SynthMorph | ||
| RegMI | ||
| PCMC | ||
| MBSS-T1 |
Clinical impact: Figure 5 presents several representative cases. The maps generated by the MBSS-T1 method demonstrate the highest quality, as evidenced by superior metrics. The rightmost column of Table 1 summarizes the clinical impact assessment results for the MBSS-T1 method, which achieved the highest quality score of . This score reflects an improvement of approximately 10% over the following best method, PCMC-T1, which had a score of . The difference in radiologist grading was statistically significant (p ). Furthermore, as demonstrated in Figure 6, the MBSS-T1 method achieves a closer fit to the expected magnetization curve (with ) and reduces artifacts, which may contribute to improved diagnostic confidence. These results suggest the MBSS-T1 method’s potential utility in clinical settings where motion artifacts can impact the accuracy of cardiac T1 mapping.
Quantitative evaluation: Table 3 summarizes our results for Exp 2 for both breath-hold (BH) and free-breathing (FB) scans. The MBSS-T1 method outperformed Synthmorph in terms of , Dice score, and Hausdorff distance. Specifically, MBSS-T1 achieved an value of for BH and for FB, representing improvements of approximately 0.1% and 0.5%, respectively, compared to Synthmorph. Additionally, MBSS-T1 produced a Dice score of for BH and for FB, reflecting improvements of approximately 3% and 4.7%, respectively, compared to Synthmorph. Furthermore, MBSS-T1 reduced the Hausdorff distance to mm for BH and mm for FB, yielding improvements of 29% and 37%, respectively, compared to Synthmorph.
Clinical impact: The rightmost column in both the breath-hold (BH) and free-breathing (FB) sections of Table 3 summarizes the clinical assessment results for the MOLLI sequence experiment. The MBSS-T1 method enhances the quality of T1 maps by effectively reducing motion artifacts, as reflected by the highest clinical scores in both BH and FB scans. Specifically, MBSS-T1 achieved mean scores of for BH and for FB scans. These scores represent improvements of approximately 20% in BH and 17% in FB compared to the next best method, Synthmorph, which had scores of and , respectively.
Method Breath Hold Free Breathing Dice HD [mm] Clinical Score Dice HD [mm] Clinical Score w.o motion correction 0.991 0.026 0.871 0.131 5.21 5.33 2.89 0.89 0.965 0.054 0.851 0.141 5.45 2.36 2.81 0.93 Siemens MyoMap N/A N/A N/A 2.84 0.66 N/A N/A N/A 3.28 0.73 SynthMorph 0.993 0.020 0.924 0.087 3.75 4.91 3.15 0.62 0.982 0.026 0.919 0.080 3.71 1.51 3.5 0.58 MBSS-T1 0.994 0.017 0.954 0.083 2.64 4.77 3.79 0.56 0.987 0.022 0.963 0.029 2.33 0.96 4.1 0.61
Figure 7 presents the reduction in motion-related distortions achieved by the MBSS-T1 method. This improved motion correction results in more precise and reliable T1 maps with fewer artifacts.
Figure 8 shows the T1 values calculated by the different methods, presented using a 16-segment AHA bullseye plot for a representative subject. For both free-breathing and breath-hold acquisitions, our MBSS-T1 method produced more homogeneous maps compared to the other methods.
5.3 Exp. 3: Consistency of cardiac T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
Figure 9 presents the ICC3 metrics for the mean myocardial T1 values across three motion correction approaches: no motion correction, Siemens MyoMap, and MBSS-T1MOLLI. Both motion correction methods, Siemens MyoMap and MBSS-T1MOLLI, demonstrated higher ICC values overall, indicating improved repeatability of the T1 maps. However, at the segmental level, neither method had a consistent advantage. In some segments, MBSS-T1 showed higher ICC values, while Siemens MyoMap performed better in others. Interestingly, for the free-breathing (FB) data, the motion correction algorithms yielded lower reproducibility than maps computed without motion correction for certain segments.
6 Discussion
This study introduced MBSS-T1, a subject-specific, self-supervised deep-learning approach for motion correction in cardiac T1 mapping. By integrating physical constraints through a signal model and anatomical constraints via segmentation-guided registration, the method addresses the challenges of motion artifacts in both free-breathing and breath-hold acquisitions. Unlike prior approaches requiring extensive training datasets, MBSS-T1 leverages a self-supervised framework, making it adaptable to diverse imaging protocols without requiring protocol-specific annotations.
The results demonstrated the effectiveness of MBSS-T1 across multiple datasets and acquisition sequences. In FB STONE acquisitions, MBSS-T1 outperformed baseline methods, achieving higher R2 values, Dice scores, and improved Hausdorff distances. For MOLLI acquisitions, the approach generated more homogeneous T1 maps, with higher clinical scores for both BH and FB acquisitions, surpassing the performance of Siemens MyoMap and baseline method without motion correction. These results highlight MBSS-T1’s robustness in mitigating motion artifacts and its potential to deliver clinically reliable T1 maps under challenging conditions.
The superior performance of MBSS-T1 compared to its preliminary version PCMC-T1, as reflected in both clinical score and segmentation metrics, can be attributed to two key factors. First, MBSS-T1 incorporates a pre-trained network specifically designed for segmentation tasks, enabling it to utilize segmentation results directly within its pipeline. This approach eliminates the need to simultaneously learn myocardium segmentation and registration, allowing the network to focus solely on the registration process and optimize the Dice loss more effectively. Second, MBSS-T1 employs a case-specific inference approach, solving the optimization problem individually for each test case rather than relying on a standard inference procedure, as is done in PCMC-T1. This tailored optimization strategy enhances robustness and accuracy, contributing to the significantly better performance of MBSS-T1.
Interestingly, the FB scans with motion-corrected data achieved superior clinical scores compared to the BH scans, which is a counterintuitive result. The effect of the breath-hold can be observed in the higher clinical scores assigned to the fitting process without motion correction, reflecting the inherent quality of the acquired data. We hypothesize that lower scores for the maps computed from BH scans using motion-correction algorithms may stem from the behavior of the motion-correction algorithms (Siemens MyoMap or MBSS-T1). For the BH scans, these algorithms likely detected minimal motion due to the breath hold, resulting in a limited capacity to correct any residual artifacts. As a result, minor artifacts may have persisted in the data. Conversely, the FB scans exhibited more substantial motion, which allowed the motion correction algorithms to address and mitigate a greater range of motion-related artifacts, ultimately improving the clinical scores.
It is important to note that the proposed method does not strictly guarantee anatomically correct deformations or a registration that fully compensates for underlying physiological motion. Instead, the method aims to align images to a reference state using physical and anatomical constraints that encourage expected signal decay behavior and foster realistic deformations. These constraints work to mitigate fitting errors that may arise due to motion, but their impact extends beyond motion-related artifacts to other inconsistencies in the data. Consequently, while the registration aligns the images in a manner suitable for the mapping network, it does not exclusively address physiological motion.
Another important consideration in this work is the ability of the methods to generate diffeomorphic deformation fields, which ensure smooth and invertible transformations. While achieving full diffeomorphism is not explicitly required in our approach, the registration backbone based on the diffeomorphic Voxelmorph framework encourages but does not guarantee diffeomorphic flows, especially in the context of cardiac imaging with complex respiratory and cardiac motion. Notably, strict diffeomorphic properties can be achieved by replacing the registration backbone in our system with any other diffeomorphic registration framework. This adaptability allows the method to be tailored for applications where diffeomorphic deformations are a critical requirement.
Further, our method is designed to operate on magnitude images and focuses specifically on aligning single-shot images acquired at different time points, addressing inter-frame motion. However, it does not account for motion occurring during the acquisition of the single-shot images (intra-frame motion). Such motion can introduce distortions that are beyond the scope of our current approach. Future work could explore the integration of motion-resilient acquisition techniques or advanced intra-frame motion modeling to overcome this limitation and further enhance the robustness of cardiac T1 mapping.
The subject-specific nature of our approach indicates that the neural network employs a Memorization Learning strategy, focusing on specific details or particular features of the individual examples rather than adopting the common Pattern Learning strategy, where deep neural networks (DNNs) learn generalized features or common patterns [36]. Future research is needed to investigate the generalization capacity of our approach, particularly when applied to larger and more diverse datasets.
While the proposed MBSS-T1 method demonstrates significant advancements in motion correction for cardiac T1 mapping, several limitations must be acknowledged to contextualize its findings and guide future improvements. First, the hyperparameters and the pre-trained segmentation network used in this study were specifically optimized for the free-breathing STONE sequence and subsequently evaluated on both the STONE free-breathing and MOLLI breath-hold and free-breathing sequences. This approach was necessitated by the limited number of MOLLI cases available in our dataset, which constrained our ability to perform robust optimization tailored specifically to MOLLI data. As a result, while our method demonstrated promising results on the MOLLI dataset, it is likely that these results could be further improved through hyperparameter optimization specifically designed for MOLLI data.
In addition, while our study demonstrates the potential of MBSS-T1 for motion-robust cardiac T1 mapping, particularly under FB conditions and suboptimal compliance with BH instructions, we acknowledge that our approach has limitations in addressing sub-optimal echo-triggering due to arrhythmias. Specifically, the assumed signal model may fail in the presence of arrhythmias due to spin history effects [6].
Further, the automated selection of the reference image based on segmentation performance may introduce registration bias, favoring the anatomical features of the chosen image. While this enhances segmentation accuracy, it may not represent the average anatomical state. Future work could explore alternatives, such as using an average template or dynamic reference selection, to reduce this potential bias.
Finally, there is room for further optimization of the network architecture to enhance its performance and generalization. One promising area for future exploration is the incorporation of different normalization layers, which have the potential to stabilize training, improve convergence, and enhance the network’s robustness.
7 Conclusion
In this study, we proposed MBSS-T1, a subject-specific, self-supervised motion correction method for robust cardiac T1 mapping. By integrating physical constraints through a signal model and anatomical constraints via segmentation, MBSS-T1 effectively mitigates motion artifacts in both free-breathing and breath-hold acquisitions. The method demonstrated superior performance compared to baseline approaches, producing more homogeneous T1 maps and achieving higher clinical scores and segmentation metrics. Importantly, the self-supervised nature of MBSS-T1 enables its adaptability to diverse imaging protocols without requiring large annotated datasets. Future work will focus on addressing current limitations, including exploring advanced motion modeling and optimization strategies, to further enhance the method’s robustness and clinical applicability.
@@ Acknowledgments
Declaration of Generative AI
During the preparation of this work, the author(s) used ChatGPT in order to improve readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.